OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=7 , (9,2) what is the slope intercept form is y=?

OpenStudy (anonymous):

y-2=7(x-9) y=7x-63+2 y=7x-61

OpenStudy (anonymous):

thank you very much mattt9 u re the best

OpenStudy (anonymous):

No problem, happy studies